OpenStudy (anonymous):

SIMPLIFY. ((Sin^4)x)/((tan^2)x)

OpenStudy (dumbcow):

\[\large \frac{\sin^{4} x}{\frac{\sin^{2}x}{\cos^{2}x}} = \sin^{2}x \cos^{2} x\]

OpenStudy (anonymous):

Thank u!